OpenStudy (anonymous):

I spent 1/4 of my money on a new book. I gave $3.50 and $4.75 on a pair of earring. When I came home I had $18.00 as left money. How much money I had in the beginning?I need help with my question. Help!

hero (hero):

gave = spent

OpenStudy (anonymous):

35 bucks

hero (hero):

x - x/4 - 3.50 - 4.75 = 18.00 x - x/4 - 8.25 = 18.00 x - x/4 = 18.00 + 8.25 x - x/4 = 26.25 4x/4 - x/4 = 26.25 (3x)/4 = 26.25 3x = 26.25 x 4 x = (26.25 x 4)/3 x = 35 You had 35.00 You spent 1/4 of that: 8.75 35 - 8.75 = 26.25 You spent 8.25 more 26.25 - 8.25 = 18.00